Okay so I'm in geometry and we're doing Segment Addition Postulate, here's the question: Find MN if N is between M and P, MN= 3x+2, NP= 18, and MP= 5x.
Answer by Theo(13342)

segment addition postulate states:

In geometry, the segment addition postulate states that, given two points A and C, a third point B lies on the line segment AC if and only if the distances between the points satisfy the equation AB + BC = AC.

change the names and it becomes:

In geometry, the segment addition postulate states that given two points M and P, a third point N lies on the line segment MP if and only if the distances between the points satisfy the equation MN + NP = MP

your overall segment is mn
n is between m and p
you are given that:
mn = 3x + 2
np = 18
mp = 5x

the addition postulates says that mn + np = mp

replace mp with 5x and mn with 3x + 2 and np with 18 and you get:

3x + 2 + 18 = 5x

combine like terms to get:

3x + 20 = 5x

subtract 3x from both sides of this equation to get:

20 = 5x - 2x

combine like terms to get:

20 = 2x

solve for x to get:

x = 10

replace x with 5 and you get:

mn = 3x + 2 = 32
np = 18
mp = 5x = 50

mn + np = mp becomes 32 + 18 = 40 which becomes 50 = 50

your solution is that mn = 32